A chemist has 35% NaCl solution and 50% NaCl solution. How many liters of each should he mix to get 10 L of NaCl Solution with 40% base concentration?
Answer by Alan3354(69443)

A chemist has 35% NaCl solution and 50% NaCl solution. How many liters of each should he mix to get 10 L of NaCl Solution with 40% base concentration?
-------------------
t = 35%
f = 50%
-----------
t + f = 10
0.35t + 0.5f = 0.4*10
t = 10 - f
0.35t + 0.5f = 0.4*10
Sub for t
0.35(10-f) + 0.5f = 4
7(10-f) + 10f = 80
70 + 3f = 80
f = 10/3 liters
t = 20/3 liters